# Unified laser stabilization and isolation on a silicon chip

### v1

@@ Abstract
Coherent optical sources serve as the backbone of optical communication, sensing and metrology, and quantum technologies [1, 2, 3]. Long coherence times and narrow linewidths allow for more precise heterodyne detection, interferometry, and probing of atomic transitions. Traditionally, the narrow linewidth optical sources required for these applications are built from table-top laser systems, but technological innovation in optical networking [4, 5], data processing [6, 7, 8], light detection and ranging (LiDAR) [9, 10, 11], and chip scale quantum computers [12, 13, 14] demand the same performance from compact and fully integrated systems. In this vein, significant efforts have been made to miniaturize and integrate the components of low noise lasers. To do this, three key components must be combined: a semiconductor gain chip or laser, an external cavity for linewidth stabilization, and an optical isolator to prevent unwanted reflections from destabilizing the laser.
@@
Recently, extremely high quality factor (Q ) silicon nitride ring cavities have been used in conjunction with chip scale semiconductor lasers to achieve Hz level linewidths on an integrated platform [20, 21], rivaling even the highest performance table-top systems. These devices work by taking advantage of the naturally occurring back-scattering in high Q resonators to provide narrow linewidth feedback that self-injection locks a semiconductor laser. As back-scattering is nearly ubiquitous in high Q resonators, this approach is reliable even though it is not deterministic. However, as we will show in this paper, when power levels in the device are high enough to access nonlinear processes in the resonator material, this technique starts to break down. This effect imposes stringent limits on the maximum operating power and quality factor of current architectures. Additionally, these devices still require external isolators.
Integrated isolators have also made significant strides recently, with demonstrations of on-chip isolators approaching the performance of stand-alone optical elements [22, 23]. However, there are significant challenges in integrating these with low noise systems. For instance, resonant devices [24, 22] would require the tuning of multiple high Q resonators to degenerate frequencies. Meanwhile, nonresonant devices can be quite large ( 1 cm) and, without a resonant filter, only provide a frequency shift of the back-reflected power reaching the laser cavity [23]. While this has been shown to prevent laser destabilization, it is not clear that it would allow for the preservation of ultra-narrow linewidth.
@@ I Theory of Operation
High quality factor ring resonators can be used to directly isolate the output of a continuous wave (CW) laser[25, 26]. This works by taking advantage of the nonreciprocity of the optical Kerr effect (third-order nonlinearity ). When one of the degenerate clockwise and counterclockwise modes of a ring is pumped, the differential action of self-phase (SPM) and cross-phase modulation (XPM) induce a differential shift in the clockwise and counterclockwise resonance frequencies. As XPM is twice as strong as SPM, the ring mode counter-propagating with respect to the pump is shifted twice as far in frequency as the pump mode, leading to the split transmission spectrum illustrated in Fig 1c. The pump can then transmit through the ring with near unity efficiency, while any power reflected back is no longer resonant with the ring, and is thus isolated.
The isolation achieved with this scheme can be found by taking the Lorentzian transmission of the detuned resonance [25, 26]:
where is the original resonance frequency of the ring, is the nonlinear refractive index, is the linear refractive index, is the loaded quality factor of the ring, is the mode volume of the ring, is the coupling efficiency of the pump to the ring, and is the input power. As the isolation ratio scales as , it is highly desirable to operate with a large quality factor (). This not only ensures the maximum possible isolation, but also reduces the power threshold required to induce isolation.
@@
(where is the Planck constant, is the laser frequency, and is the laser power) is dependent on Q, the frequency noise of a laser can be reduced by the introduction of a high Q feedback by the noise reduction factor (NRF)
where is the fraction of power fed back to the laser. As semiconductor chip lasers typically have a low quality factors () to maximize their output power, we can use the high quality factor rings () to reduce their frequency noise by orders of magnitude. In the remainder of the paper, we use Q to denote the loaded cavity Q factor (labeled as above).
To achieve simultaneous isolation and feedback stabilization, we propose the topology illustrated in Fig 1a, which we term a unified laser stabilizer. A chip laser is coupled to an isolator ring, which transmits power in the clockwise mode to an output waveguide. This output waveguide is tapped off with a directional coupler, and a fraction of the power is fed back to the laser. To reach the laser, the feedback power has to pass by the ring. As it is now traveling in the opposite direction, it interacts only with the detuned counterclockwise mode, preventing power coupling to the ring and allowing full transmission back to the laser. Meanwhile, any back-reflected power is off-resonant with the ring and gets dumped, preventing it from reaching the laser diode (Fig 1b). Here the ring acts effectively as a circulator. If the ring is critically coupled, transmission is only allowed from ports (as labeled in Fig. 1c). If the ring is not critically coupled, the allowed transmission is . The circulator behavior allows a large fraction of the output power to be fed back to the laser diode, leading to a strong reduction in the laser linewidth (Fig 1d).
@@
First, due to their different scaling, the isolation and noise reduction have different optimal loading of the ring. To maximize isolation, one should maximize , where and are the coupling rates of the ring to the two waveguides, and is the intrinsic loss rate of the ring. For , this leads to an optimum of , Fig 2a. To maximize NRF, one should maximize , which with leads to an optimum of , Fig 2b, approximately 5 times more loaded than the isolation optimum. Outside of this range ( the isolation and noise reduction factors reduce together, Fig 2c.
@@
As every system has different noise, power, and isolation requirements, the optimal device configuration will likely be significantly different in every scenario. In this paper, we chose to load the devices to (demarcated in Fig 2) and feed back 50% of the out-coupled power to demonstrate a relatively low power device with good isolation and NRF.
@@ II High-Q Stability
In addition to providing simultaneous isolation and noise reduction, the unified laser stabilizer topology solves a critical issue for self-injection locked lasers. Current schemes for the self-injection locking of lasers rely on either parasitic back-reflection in rings [29, 30, 20, 21, 31, 32], or a reflector placed after port 2 of the ring (drop port) [33, 34], Fig 3a. While these have been quite effective in the low power or low Q regimes, these schemes rely on the coupling and degeneracy of the clockwise and counterclockwise modes of the ring. If the power or Q was increased enough to break this degeneracy and provide isolation, the feedback strength would be heavily attenuated. As the noise reduction factor is proportional to the square of the feedback strength, this attenuated power would lead to a dramatic reduction in efficacy of these systems. Using published device parameters, we calculate at what power level and Q factor these state-of-the-art injection-locked lasers [20, 29, 31, 30, 35, 34, 36, 37, 38, 39, 33] will begin to fail (Supplemental Information Section 2). In many cases these failure points are within an order of magnitude of the current operating point.
In Fig 3b, we show the theoretical and experimental power-dependent back-reflection of a ring. Intrinsic back-scattering in the ring couples the clockwise and counterclockwise modes, causing a fraction of the input power to reflect. At low input power, the backreflected power increases linearly with the input power. As the input power increases, however, the clockwise and counterclockwise modes become increasingly detuned, and therfore less power is coupled backwards. This reduction is equal to the isolation ratio of the ring, here reaching a maximum of 15 dB.
By instead providing nonreciprocal feedback through the ring, Fig 3c, we can achieve a nearly linear response even at high input powers. In Fig 3d, we show the backreflected power for a device with this new topology that also shows 15 dB of isolation. The reduced input power where this is achieved is due to a higher Q factor. At low input power, there are three competing effects - the intrinsic ring back-scattering, the feedback through the ring, and the filtering the ring does to this feedback. At high power, however, the intrinsic back-scattering and feedback filtering are attenuated by the isolation ratio, allowing for a linear response (Supplemental Information Section 3). We compare the measured back-scattering ratio and the corresponding theoretical noise reduction factors of these topologies in Fig 3e-f. With current back-scattering based schemes, the presence of 15 dB isolation drops the noise reduction factor by 30 dB, destroying its stabilization capability. In contrast, the unified laser stabilization topology preserves the high noise reduction factor across all power levels.
@@ III Device Integration and Measurement
We implement the ULS using a silicon nitride on insulator platform. While these devices are currently fabricated in-house, ultra high-Q silicon nitride photonics is fully compatible with foundry CMOS processes, as demonstrated by [40, 32], making scalable integration of these devices practical. To maintain a normal dispersion profile and prevent parasitic nonlinear conversion processes, we use a 310 nm waveguide height controlled by low-pressure chemical vapor deposition (LPCVD) of silicon nitride. To enable high quality factors, we use multi-mode rings (4 m width) and clad the devices with a thick oxide layer through plasma-enhanced chemical vapor deposition (PECVD).
We first measure the isolation of the ULS with a tunable external cavity laser (ECL), Fig 4a. At each input power, we tune the laser across the ring resonance and record the backward transmission at the resonance peak. As expected, the isolation ratio follows a Lorentzian power-dependent curve proportional to the nonlinear detuning. We then edge-couple the ULS device to a semiconductor distributed feedback (DFB) laser that provides a coupled on-chip power of approximately 33 mW, Fig 4b. With a correlated self-heterodyne setup as described in [41], we measure the single side band (SSB) frequency noise power spectral density, Fig 4c. Compared to the free-running DFB laser, our hybrid device displays a noise reduction of 25-35 dB, limited at high offset frequency by the thermorefractive noise of the ring (Supplemental Information Section 5). At high frequency offset, close to the white noise floor of the laser, the SSB frequency noise power spectral density is reduced from 8200 Hz2/Hz to 14 Hz2/Hz, corresponding to instantaneous linewidths of 51 kHz and 89 Hz respectively.
@@
In this system, the combination of thermal locking and self-injection locking allows for laser stability without external feedback loops. In Fig 4f, we demonstrate the robustness of this type of laser system: when the laser is locked, it can be fully turned on and off over the timescale of at least tens of seconds. Without the self-injection locking feedback, the ring would need to be thermally tuned each time upon startup in order to align its transmission spectrum with the laser. Finally, we ensure that this topology is still capable of taking advantage of the rich dynamics that traditionally self-injection locked lasers exhibit [42, 20, 37]. By tuning the coupling gap and thus feedback phase with a piezoelectric positioner, we can observe both CW and frequency comb operation, and even second harmonic generation (Supplemental Information Sections 7-8) [39].
@@ IV Conclusion
Here we have demonstrated the first fully on-chip narrow-linewidth laser system, complete with built-in optical isolation. To do this, we combined self-injection locking and isolation in a single CMOS-compatible silicon nitride unified laser stabilizer. While previous self-injection schemes function well at low power, we show that they fail to function at higher powers. We propose a device topology that provides strong feedback independent of input power while simultaneously providing isolation. Our fabricated devices passively isolate an integrated DFB laser by 14 dB and at the same time reduce its frequency noise by 25-35 dB, and operate with turn-key reliability. By increasing the device quality factor through commercial scale fabrication [40], the isolation and noise reduction can be even further enhanced, and the power threshold for operation can be reduced (Fig 2, Supplemental Information Section 9). As many hybrid and heterogeneously integrated photonics already include low-loss silicon nitride or similar materials, these devices can be readily integrated into state-of-the-art systems for sensing, metrology, and quantum technologies.
@@ Methods
Silicon nitrite thin films (310nm) were deposited on silicon dioxide/silicon wafers with low-pressure chemical vapor deposition. Devices were then patterned with ZEP520A resist and e-beam lithography (JEOL JBX-6300FS). After development, the patterns were transferred to the silicon nitride using inductively coupled plasma etching with CHF3/CF4 chemistry. Devices were then cleaned with a Piranha solution, annealed at 1,100∘ C in a nitrogen environment, and oxide clad using spun hydrogen silsesquioxane (HSQ) and thermal oxide. The devices were annealed a final time under the same conditions and laser stealth diced to create clean facets.
@@
To verify whether the frequency noise of our device reached the thermal limit, we closely followed the work by [43, 44] to simulate thermorefractive noise (TRN) and corresponding resonance frequency fluctuations in the ring resonator using COMSOL Multiphysics. To find the thermal noise, we made use of the fluctuation-dissipation theorem (FDT), which describes the relation between fluctuations and dissipation in a thermodynamic system. In the simulation, we calculated the fundamental eigenmode of the waveguide and added a harmonic perturbative heat source in the system with the same spatial distribution as the eigenmode. We then solved the heat transfer equation in the frequency domain, calculated the dissipated heat energy during one cycle, and used FDT to convert dissipated heat energy to frequency fluctuations. In the simulations, we assumed Si3N4 waveguide () surrounded by SiO2 cladding ( on top, on bottom) with air above and Si substrate below. The material properties used in the simulation and a more detailed analysis of TRN (temperature dependency, effects of device geometry) can be found in Supplemental Information Section 1.

### v2

@@ Abstract
Coherent optical sources serve as the backbone of optical communication, sensing and metrology, and quantum technologies [1, 2, 3]. Long coherence times and narrow linewidths allow for more precise heterodyne detection, interferometry, and probing of atomic transitions. Traditionally, the narrow linewidth optical sources required for these applications are built from table-top laser systems, but technological innovation in optical networking [4, 5], data processing [6, 7, 8], light detection and ranging (LiDAR) [9, 10, 11], and chip scale quantum computers [12, 13, 14] demand the same performance from compact and fully integrated systems. In this vein, significant efforts have been made to miniaturize and integrate the components of low-noise lasers. To do this, three key components must be combined: a semiconductor gain chip or laser, an external cavity for linewidth stabilization, and an optical isolator to prevent unwanted reflections from destabilizing the laser.
@@
Recently, extremely high quality factor (Q ) silicon nitride ring cavities have been used in conjunction with chip-scale semiconductor lasers to achieve Hz level linewidths on an integrated platform [20, 21], rivaling even the highest performance table-top systems. These devices work by taking advantage of the naturally occurring back-scattering in high Q resonators to provide narrow linewidth feedback that self-injection locks a semiconductor laser. As back-scattering is nearly ubiquitous in high Q resonators, this approach is reliable even though it is not deterministic. However, as we will show in this paper, when power levels in the device are high enough to access nonlinear processes in the resonator material, this technique starts to break down. This effect imposes stringent limits on the maximum operating power and quality factor of current architectures. Additionally, these devices still require external isolators.
Integrated isolators have also made significant strides recently, with demonstrations of on-chip isolators approaching the performance of stand-alone optical elements [22, 23]. However, there are significant challenges in integrating these with low-noise systems. For instance, resonant devices [24, 22] would require the tuning of multiple high Q resonators to degenerate frequencies. Meanwhile, nonresonant devices can be quite large ( 1 cm) and, without a resonant filter, only provide a frequency shift of the back-reflected power reaching the laser cavity [23]. While this has been shown to prevent laser destabilization, it is not clear that it would allow for the preservation of ultra-narrow linewidth.
@@ I Theory of Operation
High quality factor ring resonators can be used to directly isolate the output of a continuous wave (CW) laser [25, 26, 27]. This works by taking advantage of the nonreciprocity of the optical Kerr effect (third-order nonlinearity ). When one of the degenerate clockwise and counterclockwise modes of a ring is pumped, the differential action of self-phase (SPM) and cross-phase modulation (XPM) induce a differential shift in the clockwise and counterclockwise resonance frequencies. As XPM is twice as strong as SPM, the ring mode counter-propagating with respect to the pump is shifted twice as far in frequency as the pump mode, leading to the split transmission spectrum illustrated in Fig 1c. The pump can then transmit through the ring with near unity efficiency, while any power reflected back is no longer resonant with the ring, and is thus isolated.
The maximum isolation achieved with this scheme can be found by taking the Lorentzian transmission of the detuned resonance [26, 27]:
where is the original resonance frequency of the ring, is the nonlinear refractive index, is the linear refractive index, is the loaded quality factor of the ring, is the mode volume of the ring, is the input power, and is the coupling efficiency of the pump to the ring. At the peak of the forward resonance, where isolation is maximized, the coupling efficiency depends on the coupling rates and the intrinsic loss rate . As the isolation ratio scales as , it is highly desirable to operate with a large quality factor (). This not only ensures the maximum possible isolation, but also reduces the power threshold required to induce isolation.
@@
(where is the Planck constant, is the laser frequency, and is the laser power) is dependent on Q, the fundamental frequency noise of a laser can be reduced by the introduction of a high Q feedback by the noise reduction factor (NRF)
where is the fraction of power fed back to the laser. As semiconductor chip lasers typically have a low quality factor () to maximize their output power, we can use the high quality factor rings () to reduce their frequency noise by orders of magnitude. In the remainder of the paper, we use Q to denote the loaded cavity Q factor (labeled as above).
To achieve simultaneous isolation and feedback stabilization, we propose the topology illustrated in Fig 1a, which we term a unified laser stabilizer. A chip laser is coupled to an isolator ring, which transmits power in the clockwise mode to an output waveguide. This output waveguide is tapped off with a directional coupler, and a fraction of the power is fed back to the laser. To reach the laser, the feedback power has to pass by the ring. As it is now traveling in the opposite direction, it interacts only with the detuned counterclockwise mode, preventing power coupling to the ring and allowing full transmission back to the laser. Meanwhile, any back-reflected power is off-resonant with the ring and gets dumped, preventing it from reaching the laser diode (Fig 1b). Here the ring acts effectively as a circulator. If the ring is critically coupled (), transmission is only allowed from ports (as labeled in Fig 1c). If the ring is not critically coupled, the allowed transmission is . The circulator behavior allows a large fraction of the output power to be fed back to the laser diode, leading to a strong reduction in the laser linewidth (Fig 1d).
@@
First, due to their different scaling, the isolation and noise reduction have different optimal loading of the ring. To maximize isolation, one should maximize , where and are the coupling rates of the ring to the two waveguides, and is the intrinsic loss rate of the ring. For , this leads to an optimum of , Fig 2a. To maximize NRF, one should maximize , which with leads to an optimum of , Fig 2b, approximately 2.5 times more loaded than the isolation optimum. Outside of this range ( the isolation and noise reduction factors reduce together, Fig 2c.
@@
As every system has different noise, power, and isolation requirements, the optimal device configuration will likely be significantly different in every scenario. In this paper, we chose to load the devices to (demarcated in Fig 2) and feed back 50% of the out-coupled power to demonstrate a relatively low power device with good isolation and high NRF.
@@ II High-Q Stability
In addition to providing simultaneous isolation and noise reduction, the unified laser stabilizer topology solves a critical issue for self-injection locked lasers. Current schemes for the self-injection locking of lasers rely on either parasitic back-reflection in rings [30, 31, 20, 21, 32, 33], or a reflector placed after port 2 of the ring (drop port) [34, 35], Fig 3a. While these have been quite effective in the low power or low Q regimes, these schemes rely on the coupling and degeneracy of the clockwise and counterclockwise modes of the ring. If the power or Q was increased enough to break this degeneracy and provide isolation, the feedback strength would be heavily attenuated. As the noise reduction factor is proportional to the feedback strength, this attenuated power would lead to a dramatic reduction in efficacy of these systems (Supplemental Information Sections 2-3). Using published device parameters, we calculate at what power level and Q factor these state-of-the-art injection-locked lasers [20, 30, 32, 31, 36, 35, 37, 38, 39, 40, 34] will begin to fail (Supplemental Information Section 4). In many cases, these failure points are within an order of magnitude of the current operating point.
In Fig 3b, we show the theoretical and experimental power-dependent back-reflection of a ring. Intrinsic back-scattering in the ring couples the clockwise and counterclockwise modes, causing a fraction of the input power to reflect. At low input power, the backreflected power increases linearly with the input power. As the input power increases, however, the clockwise and counterclockwise modes become increasingly detuned, and therefore less power is coupled backwards. This reduction is equal to the isolation ratio of the ring, here reaching a maximum of 15 dB.
By instead providing nonreciprocal feedback through the ring, Fig 3c, we can achieve a nearly linear response even at high input powers. In Fig 3d, we show the backreflected power for a device with this new topology that also shows 15 dB of isolation. The reduced input power where this is achieved is due to a higher Q factor. At low input power, there are three competing effects – the intrinsic ring back-scattering, the feedback through the ring, and the filtering the ring does to this feedback. At high power, however, the intrinsic back-scattering and feedback filtering are attenuated by the isolation ratio, allowing for a linear response (Supplemental Information Section 5). We compare the measured back-scattering ratio and the corresponding theoretical noise reduction factors of these topologies in Fig 3e-f. With current back-scattering based schemes, the presence of isolation directly degrades the noise reduction factor and locking range [41] (proportional to the isolation ratio and square root of the isolation ratio respectively), destroying its stabilization capability. In contrast, the unified laser stabilization topology preserves the high noise reduction factor across all power levels.
@@ III Device Integration and Measurement
We implement the ULS using a silicon nitride on insulator platform. While these devices are currently fabricated in-house, ultra high-Q silicon nitride photonics is fully compatible with foundry CMOS processes, as demonstrated by [42, 33], making scalable integration of these devices practical. To maintain a normal dispersion profile and prevent parasitic nonlinear conversion processes, we use a 310 nm waveguide height controlled by low-pressure chemical vapor deposition (LPCVD) of silicon nitride. To enable high quality factors (Supplemental Information Section 6), we use multi-mode rings ( width) and clad the devices with a thick oxide layer through plasma-enhanced chemical vapor deposition (PECVD).
We first measure the isolation of the ULS with a tunable external cavity laser (ECL), Fig 4a. At each input power, we tune the laser across the ring resonance and record the backward transmission at the resonance peak. As expected, the isolation ratio follows a Lorentzian power-dependent curve proportional to the nonlinear detuning. We then edge-couple the ULS device to a semiconductor distributed feedback (DFB) laser that provides a coupled on-chip power of approximately 33 mW, Fig 4b. With a correlated self-heterodyne setup as described in [43], we measure the single sideband (SSB) frequency noise power spectral density, Fig 4c. Compared to the free-running DFB laser, our hybrid device displays a noise reduction of 25-35 dB, limited at high offset frequency by the thermorefractive noise of the ring. At high frequency offset, close to the white noise floor of the laser, the SSB frequency noise power spectral density is reduced from 8200 Hz2/Hz to 14 Hz2/Hz, corresponding to instantaneous linewidths of 51 kHz and 89 Hz respectively (Supplemental Information Sections 7-9).
@@
In this system, the combination of thermal locking and self-injection locking allows for laser stability without external feedback loops. In Fig 4f, we demonstrate the robustness of this type of laser system: when the laser is locked, it can be fully turned on and off over the timescale of at least tens of seconds. Without the self-injection locking feedback, the ring would need to be thermally tuned each time upon startup in order to align its transmission spectrum with the laser. Finally, we ensure that this topology is still capable of taking advantage of the rich dynamics that traditionally self-injection locked lasers exhibit [44, 20, 38, 32]. The control of feedback phase is a critical feature to enable or to avoid the comb formation when self-injection locked [32] (Supplemental Information Section 11). By tuning the coupling gap and thus feedback phase with a piezoelectric positioner, we can observe both CW and frequency comb operation, and even second harmonic generation (Supplemental Information Sections 11-12) [40].
@@ IV Conclusion
Here we have demonstrated the first fully on-chip narrow-linewidth laser system, complete with built-in optical isolation. To do this, we combined self-injection locking and isolation in a single CMOS-compatible silicon nitride unified laser stabilizer. While previous self-injection schemes function well at low power, we show that they fail to function at higher powers. We propose a device topology that provides strong feedback independent of input power while simultaneously providing isolation. Our fabricated devices passively isolate an integrated DFB laser by 14 dB and at the same time reduce its frequency noise by 25-35 dB, and operate with turn-key reliability. By increasing the device quality factor through commercial scale fabrication [42], the isolation and noise reduction can be even further enhanced, and the power threshold for operation can be reduced (Fig 2 and Supplemental Information Section 13). While we demonstrated a ULS that stabilizes a laser through injection locking, a ULS could instead be used as one of the mirrors inside a laser cavity itself. This would lead to a lower insertion loss, enabling even higher isolation and narrower linewidths. As many hybrid and heterogeneously integrated photonics already include low-loss silicon nitride or similar materials, these devices can be readily integrated into state-of-the-art systems for sensing, metrology, and quantum technologies.
@@ Methods
Silicon nitride thin films (310nm) were deposited on silicon dioxide/silicon wafers with low-pressure chemical vapor deposition. Devices were then patterned with ZEP520A resist and e-beam lithography (JEOL JBX-6300FS). After development, the patterns were transferred to the silicon nitride using inductively coupled plasma etching with CHF3/CF4 chemistry. Devices were then cleaned with a Piranha solution, annealed at 1,100∘ C in a nitrogen environment, and oxide clad using spun hydrogen silsesquioxane (HSQ) and thermal oxide. The devices were annealed a final time under the same conditions and laser stealth diced to create clean facets.
@@
To verify whether the frequency noise of our device reached the thermal limit, we closely followed the work by [45, 46] to simulate thermorefractive noise (TRN) and corresponding resonance frequency fluctuations in the ring resonator. Thermal noise was found using fluctuation-dissipation theorem (FDT) and finite-element method (COMSOL Multiphysics). Specifically, we added a harmonic perturbative heat source in the system with the same spatial distribution as the fundamental waveguide eigenmode, solved the heat transfer equation in the frequency domain, calculated the dissipated heat energy during one cycle, and used FDT to convert dissipated heat energy to frequency fluctuations. In the simulations, we assumed Si3N4 waveguide () surrounded by SiO2 cladding ( on top, on bottom) with air above and Si substrate below. The material properties used in the simulation and a more detailed analysis of TRN (temperature dependency, effects of device geometry) are described in Supplemental Information Section 1.
Data Availability
All data are available from the corresponding authors upon reasonable request.
Additional Information
Supplementary Information is available for this paper. Correspondence and requests for materials should be addressed to adwhite@stanford.edu, gahn@stanford.edu.